# Safe MPC Alignment with Human Directional Feedback

### v2

@@ B-D1 Environment
The arm starts at a configuration in which the end-effector position $\boldsymbol{r}_{\text{init}}=[0.22,-0.5,0.48]^{\mkern-1.5mu\mathsf{T}}$ and altitude $\boldsymbol{q}_{\text{init}}=[0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$. The reaching target position $\boldsymbol{r}^{*}$ will be randomly chosen from three options $[-0.6,-0.5,0.4]^{\mkern-1.5mu\mathsf{T}}$, $[-0.6,-0.5,0.5]^{\mkern-1.5mu\mathsf{T}}$, $[-0.6,-0.5,0.6]^{\mkern-1.5mu\mathsf{T}}$. The reaching target attitude is $\boldsymbol{q}^{*}=[0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$. The slot size and position are unknown except its rough x-position which is $x_{\text{slot}}\approx-0.15$.
@@
which is used to constrain the $\mathrm{RBF}_{i}$ only taking effect around $x$ range $[x_{\text{slot}}-\frac{w}{2},x_{\text{slot}}+\frac{w}{2}]$ (recall slot’s $x$ location, slot, is a approximate value). In the above definition, $w=0.3$, $k_{z}=0.05$, $\epsilon_{z}=12$ and $\epsilon_{q}=1.8$. The center of positional $\mathrm{RBF}$, ${c_{z,i}}$, is evenly chosen on interval $[0,1]$, and the center of orientation $\mathrm{RBF}$, ${\boldsymbol{q}_{c,i}}$, is generated by $(\cos{\frac{\alpha}{2}},\sin{\frac{\alpha}{2}},0,0)*\boldsymbol{q}_{\text{init}}$ with $\alpha$ evenly choosen from $[-\pi/4,3\pi/4]$ to make sure ${\boldsymbol{q}_{c,i}}$ covers all possible orientations. $\beta_{t}=0.9^{t}$ if $t<5$ and 00 otherwise. The bounds of the initial hypothesis space $\Theta_{0}$ is set to $\boldsymbol{c}_{l}=[-3,\dots,-3]^{\mkern-1.5mu\mathsf{T}}$ and $\boldsymbol{c}_{u}=[5,\dots,5]^{\mkern-1.5mu\mathsf{T}}$. $\gamma$ is set to 1.5. This initialization leads to a conservative MPC policy, preventing the arm from approaching the target at the beginning.
@@ B-D2 Human Correction Interface
The motion of the robot arm is rendered in the free-view mode in MuJoCo visualization (Fig. 16), and users can drag to select his/her comfortable view. The slot and the random reaching target are visible to the users.
Human users apply their corrections to the robot arm via a keyboard. Users are informed with the following key mapping between the key press and the corresponding correction in the game. The safety constraint (60) in Franka arm’s safe MPC policy is updated immediately once a user correction is detected. When user press "ENTER", the arm is reset to the initial position and the reaching target will also change.
TABLE VI: keyboard commands in robot arm user study
Key Correction Interpretation
"left" $[0,0,0,0,0,1]^{\mkern-1.5mu\mathsf{T}}$ Clock-wise rotation of the gripper
"right" $[0,0,0,0,0,-1]^{\mkern-1.5mu\mathsf{T}}$ Anti-clock-wise rotation of the gripper
"shift" $[0,0,0,-1,0,0]^{\mkern-1.5mu\mathsf{T}}$ Go up in world frame
"ctrl" $[0,0,0,1,0,0]^{\mkern-1.5mu\mathsf{T}}$ Go down in world frame
"Enter" N/A Reset robot position and switch target.

### v3

@@ B-D1 Environment
The arm starts at a configuration in which the end-effector position $\boldsymbol{r}_{\text{init}}=[0.22,-0.5,0.48]^{\mkern-1.5mu\mathsf{T}}$ and altitude $\boldsymbol{q}_{\text{init}}=[0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$. The reaching target position $\boldsymbol{r}^{*}$ will be randomly chosen from three options $0.6,-0.5,0.4]^{\mkern-1.5mu\mathsf{T}}$, $0.6,-0.5,0.5]^{\mkern-1.5mu\mathsf{T}}$, $0.6,-0.5,0.6]^{\mkern-1.5mu\mathsf{T}}$. The reaching target attitude is $\boldsymbol{q}^{*}=[0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$. The slot size and position are unknown except its rough x-position which is $x_{\text{slot}}\approx-0.15$.
@@
TABLE VI: keyboard commands in robot arm user study
Key Correction Interpretation
"left" $0,0,0,0,0,1]^{\mkern-1.5mu\mathsf{T}}$ Clock-wise rotation of the gripper
"right" $0,0,0,0,0,-1]^{\mkern-1.5mu\mathsf{T}}$ Anti-clock-wise rotation of the gripper
"shift" $0,0,0,-1,0,0]^{\mkern-1.5mu\mathsf{T}}$ Go up in world frame
"ctrl" $0,0,0,1,0,0]^{\mkern-1.5mu\mathsf{T}}$ Go down in world frame
"Enter" N/A Reset robot position and switch target.
@@
which is used to constrain the $\mathrm{RBF}_{i}$ only taking effect around $x$ range $x_{\text{slot}}-\frac{w}{2},x_{\text{slot}}+\frac{w}{2}]$ (recall slot’s $x$ location, slot, is a approximate value). In the above definition, $w=0.3$, $k_{z}=0.05$, $\epsilon_{z}=12$ and $\epsilon_{q}=1.8$. The center of positional $\mathrm{RBF}$, $c_{z,i}$}, is evenly chosen on interval $[0,1]$, and the center of orientation $\mathrm{RBF}$, $\boldsymbol{q}_{c,i}$}, is generated by $(\cos{\frac{\alpha}{2}},\sin{\frac{\alpha}{2}},0,0)*\boldsymbol{q}_{\text{init}}$ with $\alpha$ evenly choosen from $\pi/4,3\pi/4]$ to make sure $\boldsymbol{q}_{c,i}$} covers all possible orientations. $\beta_{t}=0.9^{t}$ if $t<5$ and 0 otherwise. The bounds of the initial hypothesis space $\Theta_{0}$ is set to $\boldsymbol{c}_{l}=[-3,\dots,-3]^{\mkern-1.5mu\mathsf{T}}$ and $\boldsymbol{c}_{u}=[5,\dots,5]^{\mkern-1.5mu\mathsf{T}}$. $\gamma$ is set to 1.5. This initialization leads to a conservative MPC policy, preventing the arm from approaching the target at the beginning.
@@ B-D2 Human Correction Interface
The robot arm’s motion is visualized in MuJoCo using a free-view mode (Fig. 16(b)), allowing users to adjust the camera to a preferred angle. The slot and the randomized reaching target remain visible throughout the task. Users provide corrections to the robot arm via keyboard input, following the specified key–command mappings. Whenever a correction is detected, the safety constraint (61) in the Franka arm’s safe MPC policy is updated immediately. Pressing ENTER resets the arm to its initial position and randomizes the reaching target.
[Uncaptioned image] Zhixian Xie Zhixian Xie is currently a Ph.D. student in Arizona State University, Tempe, AZ, USA. In 2022, he received B. Eng degree in automation at Shanghai Jiaotong University, Shanghai, China. His research focuses mainly on controls, robotic learning from human and robot dexterous manipulation.
[Uncaptioned image] Wenlong Zhang Wenlong Zhang is an associate professor in the School of Manufacturing Systems and Networks at Arizona State University, where he serves as the Research Director. He is a core faculty member in the Robotics and Autonomous Systems and Systems Engineering PhD programs. He joined ASU faculty in 2015 after earning his doctoral degree in mechanical engineering from the University of California, Berkeley. At ASU, he leads the Robotics and Intelligent Systems Laboratory (RISE Lab), and teaches undergraduate- and graduate-level classes on dynamics, control systems, optimization, and robotics.
[Uncaptioned image] Yi Ren Yi (Max) Ren is an associate professor of aerospace and mechanical engineering in the School for Engineering of Matter, Transport and Energy at Arizona State University. he directs the Design Informatics Laboratory at ASU. His group studies computational and data-driven methods that will augment or transform the activities of engineering and industrial design. Ren’s current research interests include optimization, product/configuration design, human-computer interaction and machine learning.
[Uncaptioned image] Zhaoran Wang Zhaoran Wang is an associate professor with Northwestern University, working at the interface of machine learning, statistics, and optimization. He is the recipient of the AISTATS (Artificial Intelligence and Statistics Conference) notable paper award, ASA (American Statistical Association) best student paper in statistical learning and data mining, INFORMS (Institute for Operations Research and the Management Sciences) best student paper finalist in data mining, Microsoft PhD Fellowship, SimonsBerkeley/J.P. Morgan AI Research Fellowship, Amazon Machine Learning Research Award, and NSF CAREER Award.
[Uncaptioned image] George J. Pappas George J. Pappas is the UPS Foundation Professor at the Department of Electrical and Systems Engineering at the University of Pennsylvania. He also holds a secondary appointment in the Departments of Computer and Information Sciences, and Mechanical Engineering and Applied Mechanics. Previously, he served as the Deputy Dean for Research in the School of Engineering and Applied Science. Pappas’s research interest focus on control systems, robotics and autonomous systems, formal methods, machine learning for safe and secure cyber-physical systems. He has received numerous awards including the NSF PECASE, the Antonio Ruberti Young Researcher Prize, the George S. Axelby Award, the O. Hugo Schuck Best Paper Award, and the George H. Heilmeier Faculty Excellence Award. Pappas has mentored more than fifty students and postdocs, now faculty in leading universities worldwide. He is a Fellow of IEEE, IFAC, and was elected to the National Academy of Engineering in 2024.
[Uncaptioned image] Wanxin Jin Wanxin Jin is an assistant professor the School for Engineering of Matter, Transport and Energy at Arizona State University (ASU). Previously, he was a postdoctoral fellow at the GRASP Laboratory at the University of Pennsylvania. He obtained his doctorate from Purdue University in 2021. His research includes robotics, control and machine learning, with a focus on the autonomy of robots interacting with humans and objects.